A portfolio has 80 shares of Stock A that sell for $32 per share and 105 shares of Stock B that sell for $29 per share.

  

Required:
(a) What is the portfolio weight of Stock A?
(Click to select)  0.4467  0.4307  0.4667  0.5446  0.4567

  

(b) What is the portfolio weight of Stock B?
(Click to select)  0.5533  0.4954  0.5341  0.5433  0.5161
Answer #1

BWNiT7b8LUwAAAABJRU5ErkJggg==

Weight of stock A =2560/5605= 0.4467

weight of stock B =3045/5605=0.5433